# A note on generalized tensor CUR approximation for tensor pairs and tensor triplets based on the tubal product

### v2

@@ Abstract
In this note, we briefly present a generalized tensor CUR (GTCUR) approximation for tensor pairs $(\underline{\bf X},\underline{\bf Y})$ and tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ based on the tubal product (t-product). We use the tensor Discrete Empirical Interpolation Method (TDEIM) to do these extensions. We show how the TDEIM can be utilized to generalize the classical tensor CUR (TCUR) approximation, which acts only on a single tensor, to jointly compute the TCUR of two and three tensors. This approach can be used to sample relevant lateral/horizontal slices of one data tensor relative to one or two other data tensors. For some special cases, the Generalized TCUR (GTCUR) approximation is reduced to the classical TCUR for both tensor pairs and tensor triplets in a similar fashion as shown for the matrices.
@@
[label1]organization=Center for Artificial Intelligence Technology, Skolkovo Institute of Science and Technology, Moscow, Russia, s.asl@skoltech.ru,
@@ 1 Introduction
where $\alpha_{i}^{2}+\beta_{i}^{2}=1$ with the ratios $\beta_{i}/\alpha_{i}$ of nondecreasing order for $i=1,2,\ldots,I_{2}$. Here, ${\bf U}\in\mathbb{R}^{I_{1}\times I_{1}},\,{\bf V}\in\mathbb{R}^{I_{3}\times I_{3}}$ are orthogonal, while ${\bf Z}\in\mathbb{R}^{I_{2}\times I_{2}}$ is a nonsingular matrix. As can be seen, the GSVD shares the common factor matrix ${\bf Z}$ between the given data matrices ${\bf X}$ and ${\bf Y}$. On the other hand, from the theory of the MCUR approximation, we know that by applying the Discrete Empirical Interpolation Method (DEIM) [6] to the basis ${\bf Z}$ (to be discussed in Section 4), we can sample appropriate column indices from the data matrices ${\bf X}$ and ${\bf Y}$. Since the same basis matrix ${\bf Z}$ is used within the DEIM, the same indices will be produced for sampling the columns of the matrix ${\bf X}$ and the matrix ${\bf Y}$. Besides, the matrices ${\bf U}$ and ${\bf V}$ can be used to sample row indices of the matrices ${\bf X}$ and ${\bf Y}$, respectively. This means that the selected row indices for these matrices can be different. As a result, for two matrices ${\bf X}$ and ${\bf Y}$ with the same number of columns, we can use the identical column indices for their MCUR approximations. The authors in [2] show that this idea can be potentially used in applications where one is interested in extracting the most discriminative features from one data set relative to another data set. It is also motivated for recovering data matrices perturbed with colored noise and subgroup discovery application. Following this idea, in [3], the authors consider triplet matrices $({\bf X},{\bf Y},{\bf Z})$ and use the restricted SVD [7] to build their common factor matrices. Then, the DEIM is applied to the common factor matrices to sample columns and rows. It is shown that this approach can be used to extract the most discriminative features from one data set relative to two other datasets. The GMCUR for tensor triplets is also motivated by the canonical correlation analysis and was successfully used for recovering datasets perturbed with colored noise, See [2, 3] for details.
@@
In this note, we focus on the tensor SVD (t-SVD) [8] that is defined based on the tubal product (t-product). The t-SVD has similar properties as the classical SVD. In particular, in contrast to the Tucker decomposition [9, 10] or Canonical polyadic decomposition [11, 12], its truncation provides the best low tubal rank approximation in the least-squares sense. The GSVD was generalized to tensors based on the t-SVD in [13]. As real-world datasets often have multidimensional structure, it is more favorable to generalize the ideas proposed in [2, 14] to tensors so that the structure of the data tensors is preserved. Indeed, we aim at achieving this target and also motivated by the works [2, 3], we extend the tensor CUR (TCUR) approximation [15] to a pair and a triplet of tensors based on the t-product. We refer to this extended TCUR approximation as the generalized TCUR (GTCUR). To sample important lateral and horizontal slices, we employ the tensor DEIM (TDEIM) index selection algorithm [16], which we recently developed and this plays a key role in the development of the GTCUR method. For a given pair or triplet of third order tensors, the TDEIM aids in computing the same lateral/horizontal slices as in the matrix case. We remark that alternative sampling procedures, such as top tubal leverage scores [17, 15] or its sampling version, may be used in place of the TDEIM, but because the TDEIM has been shown to give the optimal sample method [6], we adopt it in our research.
@@
The remainder of this note is organized as follows. We first present some basic tensor concepts in Section 2. The t-SVD is introduced in Section 2.1 with some intuitions behind this model. The Tensor Discrete Empirical Interpolation Method (TDEIM) as an extension of the DEIM method is presented in Section 4. The matrix and tensor cross approximations are outlined in Section 3. In Sections 5 and 6, we show how the matrix cross approximation for matrix pairs/triplets can be generalized to tensor pairs and tensor triplets, respectively. The conclusion is given in Section 7.
@@ 2 Preliminaries
[…]
@@
(Moore-Penrose pseudoinverse of a tensor) Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ be given. The Moore-Penrose (MP) pseudoinverse of the tensor $\underline{\bf X}$ is denoted by $\underline{\bf X}^{{\dagger}}\in\mathbb{R}^{I_{2}\times I_{1}\times I_{3}}$ is a unique tensor satisfying the following four equations:
@@
The inverse of a tensor is a special case of the Moore-Penrose pseudoinverse of tensors. The inverse of $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}}$ is denoted by $\underline{\bf X}^{-1}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}}$ is a unique tensor satisfying $\underline{\bf X}*\underline{\bf X}^{-1}=\underline{\bf X}^{-1}*\underline{\bf X}=\underline{\bf I},$ where $\underline{\bf I}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}}$ is the identity tensor The inverse of a tensor can be also computed in the Fourier domain by replacing the MATLAB command “inv” in Line 3 of Algorithm 2.
@@ 2 $\widehat{\underline{\mathbf{Y}}}={\rm fft}\left({\underline{\mathbf{Y}},[],3}\right)$;
Output : Moore-Penrose $\underline{\mathbf{X}}^{{\dagger}}\in\mathbb{R}^{I_{2}\times I_{1}\times I_{3}}$
@@ 3       $\widehat{\underline{\mathbf{C}}}\left({:,:,i}\right)={\rm pinv}\,(\widehat{{\bf X}}(:,:,i))$;
Algorithm 2 Fast Moore-Penrose computation of the tensor $\underline{\bf X}$
@@ 3 Matrix CUR (MCUR) approximation of a single matrix and matrix pairs
The MCUR approximation of matrix pairs (matrices with the same number of columns) as the extension of the classical MCUR approximations were presented in [2]. The key idea of this generalization is to select columns from two matrices that share the same column indices in order to extract the features of one dataset that are most pertinent to the other. Here, the common factor matrix produced by the GSVD of the two matrices is used to sample the columns and rows using the Discrete Empirical Interpolation Method (DEIM) technique. The MCUR approximation of triplet matrices was also presented in [3] , with the goal of identifying pertinent characteristics of one data matrix relative to two other matrices. The authors in this instance apply the RSVD to produce two common factor matrices, to which the DEIM is then applied to sample the column and row indices. We first discuss the DEIM extension to tensors based on the t-product in the next section before using it to extend the TCUR technique to tensor pairs and tensor triplets in Sections 5 and 6, respectively.
@@ 4 Tensor Discrete Empirical Interpolation Method (TDEIM) for lateral/horizontal slice sampling
We have recently generalized the DEIM method to the tensor case based on the t-product [16] that we call it tensor DEIM (TDEIM). It has been experimentally shown that the TDEIM achieved the best sampling accuracy compared to the existing sampling techniques such as the top tubal leverage score sampling and uniform sampling. The TDEIM to select horizontal slices is presented in Algorithm 4. A key concept in the TDEIM is the interpolatory projector, as we now define it. For a given set of indices ${\bf s}\in\mathbb{R}^{R}$, the full tubal-rank tensor [^1] A tensor with linear independent lateral slices, for example, the tensor $\underline{\bf U}$ obtained from the t-SVD can be used. $\underline{\bf U}\in\mathbb{R}^{I_{1}\times R\times I_{3}}$ and considering $\underline{\bf S}=\underline{\bf I}(:,{\bf s},:)\in\mathbb{R}^{I_{1}\times R\times I_{3}}$, as the selection tensor where $\underline{\bf I}\in\mathbb{R}^{R\times R\times I_{3}}$ is an identity tensor, we build the tensorial oblique projection operator defined as follows
@@
$s_{j}=\arg\max_{1\leq i\leq I_{1}}\|\underline{\bf R}(i,j,:)\|_{2},$
is set as our new index $s_{j}$. This will be the new sampled horizontal slice’s index. The same procedure can be used to select lateral slices where the basis tensor $\underline{\bf V}$ obtained from the t-SVD of the tensor $\underline{\bf X}$ should be used.
@@
that will be used in the next theorem. Here, $\underline{\bf V}\in\mathbb{R}^{I_{2}\times R\times I_{3}}$ is a basis tensor for the subspace of lateral slices of the original data tensor, $\underline{\bf Q}=\underline{\bf I}(:,{\bf s}_{j-1},:)$ is a tensor of some sampled lateral slices of the identity tensor specified by the index set ${\bf s}_{j-1}=\{s_{1},s_{2},\ldots,s_{j-1}\},\,\widehat{\underline{\bf Q}}={\rm fft}(\underline{\bf Q},[],3),\,\widehat{\underline{\bf V}}={\rm fft}(\underline{\bf V},[],3),\,\widehat{\underline{\bf Q}}_{i}=\widehat{\underline{\bf Q}}(:,:,i),$ and $\widehat{\underline{\bf V}}_{i}=\widehat{\underline{\bf V}}(:,:,i)$.
The next theorem demonstrates the upper error bound of the TCUR approximation obtained by the TDEIM sampling method.
@@
[16] Suppose $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $1\leq R<\min(I_{1},I_{2})$. Assume that horizontal slice and lateral slice indices ${\bf p}$ and ${\bf q}$ give full tubal rank tensors $\underline{\bf C}=\underline{\bf X}(:,{\bf q},:)=\underline{\bf X}*\underline{\mathcal{Q}}$ and $\underline{\bf R}=\underline{\bf X}({\bf p},:,:)={\underline{\mathcal{P}}}*\underline{\bf X}$, where $\underline{\mathcal{P}}$ and $\underline{\mathcal{Q}}$ are tensorial interpolation projectors for horizontal and lateral slice sampling [^2] The tensorial interpolation projector for the lateral slice sampling is defined as $\underline{\mathcal{Q}}=\underline{{\bf V}}*(\underline{{\bf Q}}^{T}*\underline{{\bf V}})^{-1}*\underline{{\bf Q}}^{T}$, where $\underline{{\bf V}}$ is a basis tensor and $\underline{{\bf Q}}$ is the selection tensor., respectively with corresponding finite error constants $\tilde{\eta}_{p},\,\tilde{\eta}_{q}$ defined in (8) and set $\underline{\bf U}=\underline{{\bf C}}^{{\dagger}}*\underline{\bf X}*\underline{\bf R}^{{\dagger}}$. Then
@@ 5 Tensor CUR (TCUR) approximation for tensor pairs based on the t-product
Inspired by the success of the MCUR approximation, several generalizations of these approaches to tensors have been proposed. For example, see [30] for the Tucker model, see [31, 32] for the TT decomposition, see [15] for the t-SVD, and see [33] for the TC/TR decomposition. In this note, we focus on the t-SVD case and extend it to a pair of tensors. The TCUR approximation was proposed in [15] for low tubal rank approximations of tensors. For a given tensor $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, the TCUR approximation is computed as follows
$\underline{\bf X}\approx\underline{\bf C}*\underline{\bf U}*\underline{\bf R},$ (10)
where $\underline{\bf C}\in\mathbb{R}^{I_{1}\times R_{1}\times I_{3}}$ and $\underline{\bf R}\in\mathbb{R}^{R_{2}\times I_{2}\times I_{3}}$ are collections of lateral and horizontal slices while the middle tensor $\underline{\bf U}\in\mathbb{R}^{R_{1}\times R_{2}\times I_{3}}$ needs to be computed so that the approximation error of (10) is as small as possible. Motivated by the GMCUR approximation developed in [2], we propose an extended version of the TCUR approximation written in (10). The TCUR approximation is visualized in Figure 1. The intuition behind this extension is similar to the matrix case. More precisely, the indices of selected lateral slices are the same for two data tensors (see Figure 2 for a graphical illustration). The formal definition of the GTCUR for tensor pairs is presented below.
Figure 1: Visualization of the tensor CUR (GTCUR) approximation for the tensor $\underline{\bf X}$ [34].
Figure 2: Visualization of the generalized tensor CUR (GTCUR) approximation for the tensor pairs $(\underline{\bf X},\underline{\bf Y})$. The indices of the selected lateral slices of the tensors $\underline{\bf X}$ and $\underline{\bf Y}$ are identical while this is not necessarily true for the horizontal slices.
@@
Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\underline{\bf Y}\in\mathbb{R}^{I_{4}\times I_{2}\times I_{3}}$ be of full tubal rank, with $I_{1}\geq I_{2}$ and $I_{4}\geq I_{2}$. A GTCUR approximation of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$ with the tubal rank $R,$ is defined as follows
@@
where $\underline{\bf S}_{1}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\underline{\bf S}_{2}\in\mathbb{R}^{I_{4}\times R\times I_{3}}$ and $\underline{\bf P}\in\mathbb{R}^{I_{2}\times R\times I_{3}}$ are the index selection tensors ($R<I_{2}$).
Assume that ${\bf s}_{1},$ and ${\bf s}_{2}$ are the horizontal slice indices of the tensors $\underline{\bf X}$ and $\underline{\bf Y}$, respectively. Also, let ${\bf p}$ be the shared lateral slice indices for both the $\underline{\bf X}$ and $\underline{\bf Y}$ tensors. Then, formulations (11)-(12) can be equivalently written as
@@
The GTCUR for tensor pairs computes the TCUR approximations of two given data tensors $\underline{\bf X}$ and $\underline{\bf Y}$ and is motivated by the generalized t-SVD (GTSVD). For given data tensors $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\underline{\bf Y}\in\mathbb{R}^{I_{4}\times I_{2}\times I_{3}}$, the GTSVD guarantees the following decompositions [13]

### v3

@@ Abstract
In this note, we briefly present a generalized tensor CUR (GTCUR) approximation for tensor pairs $(\underline{\bf X},\underline{\bf Y})$ and tensor triplets $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ based on the tubal product (t-product). We use the tensor Discrete Empirical Interpolation Method (TDEIM) to do these extensions. We show how the TDEIM can be utilized to generalize the classical tensor CUR (TCUR) approximation, which acts only on a single tensor, to jointly compute the TCUR of two and three tensors. This approach can be used to sample relevant lateral/horizontal slices of one data tensor relative to one or two other data tensors. For some special cases, the Generalized TCUR (GTCUR) method is reduced to the classical TCUR for both tensor pairs and tensor triplets in a similar fashion as shown for the matrices.
@@
[…]
@@ 1 Introduction
where $\alpha_{i}^{2}+\beta_{i}^{2}=1$ with the ratios $\beta_{i}/\alpha_{i}$ of nondecreasing order for $i=1,2,\ldots,I_{2}$. Here, ${\bf U}\in\mathbb{R}^{I_{1}\times I_{1}},\,{\bf V}\in\mathbb{R}^{I_{3}\times I_{3}}$ are orthogonal, while ${\bf Z}\in\mathbb{R}^{I_{2}\times I_{2}}$ is a nonsingular matrix. As can be seen, the GSVD shares the common factor matrix ${\bf Z}$ between the given data matrices ${\bf X}$ and ${\bf Y}$. On the other hand, from the theory of the MCUR approximation, we know that by applying the Discrete Empirical Interpolation Method (DEIM) [6] to the basis ${\bf Z}$ (to be discussed in Section 3), we can sample appropriate column indices from the data matrices ${\bf X}$ and ${\bf Y}$. Since the same basis matrix ${\bf Z}$ is used within the DEIM, the same indices will be produced for sampling the columns of the matrix ${\bf X}$ and the matrix ${\bf Y}$. Besides, the matrices ${\bf U}$ and ${\bf V}$ can be used to sample row indices of the matrices ${\bf X}$ and ${\bf Y}$, respectively. This means that the selected row indices for these matrices can be different. As a result, for two matrices ${\bf X}$ and ${\bf Y}$ with the same number of columns, we can use the identical column indices for their MCUR approximations. The authors in [2] show that this idea can be potentially used in applications where one is interested in extracting the most discriminative features from one data set relative to another data set. It is also motivated for recovering data matrices perturbed with colored noise and subgroup discovery application. Following this idea, in [3], the authors consider triplet tensors $(\underline{\bf X},\underline{\bf Y},\underline{\bf Z})$ and use the restricted SVD [7] to build their common factor matrices. Then, the DEIM is applied to the common factor matrices to sample columns and rows. It is shown that this approach can be used to extract the most discriminative features from one data set relative to another data set. The GMCUR for tensor triplets is also motivated by the canonical correlation analysis and was successfully used for recovering datasets perturbed with colored noise, See [2, 2] for details.
@@
In this note, due to the applications of tensors [8, 9, 10], we focus on the tensor SVD (t-SVD) [11] that is defined based on the tubal product (t-product). The t-SVD has similar properties as the classical SVD. In particular, in contrast to the Tucker decomposition [12, 13] or Canonical polyadic decomposition [14, 15], its truncation provides the best low tubal rank approximation in the least-squares sense. The GSVD was generalized to tensors based on the t-SVD in [16, 17]. Motivated by the works [2, 3], we extend the tensor CUR (TCUR) approximation [18] to a pair and a triplet of tensors based on the t-product and we refer to as the generalized TCUR (GTCUR). As real-world datasets often have multidimensional structure, it is more favorable to generalize the ideas proposed in [2, 19] to tensors so that the structure of the datatensors are preserved. Indeed, we aim at achieving this target and present several generalizations of GMCUR to tensors. To this end, we use the tensor DEIM (TDEIM) index selection algorithm [20], which we recently developed to sample important lateral and horizontal slices, and this plays a key role in the development of the GTCUR method. Similar to the matrix case, the TDEIM helps to compute the same lateral/horizontal slices for a given pair or triplet of third order tensors. We remark that alternative sampling procedures, such as top tubal leverage scores [21, 18] or its sampling version, may be employed in place of the TDEIM, but because the TDEIM has been shown to give the optimal sample method [6], we adopt it in our research.
@@
The remainder of this note is organized as follows. We first present some basic tensor concepts in Section 2. The t-SVD is introduced in Section 2.1 with some intuitions behind this model. The matrix cross approximation is outlined in Section 3 and the Tensor Discrete Empirical Interpolation Method (TDEIM) as an extension of the DEIM method is presented in Section 4. In Sections 5 and 6, we show how the matrix cross approximation for matrix pairs/triplets can be generalized to tensor pairs and tensor triplets, respectively. The conclusion is given in Section 7.
@@ 2 Preliminaries
Before presenting the main materials, we first introduce the basic notations and definitions. An underlined bold capital case letter, a bold capital case letter, and a bold lower case letter are used to represent a tensor, a matrix, and a vector, respectively. For a third-order tensor $\underline{{\bf X}},$ we call the slices $\underline{{\bf X}}(:,:,k),\,\underline{{\bf X}}(:,j,:),\,\underline{{\bf X}}(i,:,:)$ frontal, lateral, and horizontal slices. For simplicity, sometimes for a frontal slice $\underline{\bf X}(:,:,i)$, we use the notation $\underline{\bf X}_{i}$. For a third-order tensor $\underline{\bf X},$ the fiber $\underline{{\bf X}}(i,j,:)$ is called a tube. The matrix’s component-wise complex conjugate is denoted by the symbol “conj”. To indicate a subset of a matrix or tensor, we use the MATLAB notations. For instance, for a given data matrix ${\bf X}$, by ${\bf X}(:,\mathcal{J})$ and ${\bf X}(\mathcal{I},:)$ we mean two matrices sampling a part of the rows and columns of the matrix ${\bf X}$, respectively, where $\mathcal{I}\subset\{1,2,\ldots,I_{2}\}$ and $\mathcal{J}\subset\{1,2,\ldots,I_{1}\}$. We need to present Definitions 1-5 to introduce the tensor SVD (t-SVD) model.
@@
(Moore–Penrose pseudoinverse of a tensor) Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ be given. The Moore-Penrose (MP) pseudoinverse of the tensor $\underline{\bf X}$ is denoted by $\underline{\bf X}^{{\dagger}}\in\mathbb{R}^{I_{2}\times I_{1}\times I_{3}}$ is a unique tensor satisfying the following four equations:
@@ 2 $\widehat{\underline{\mathbf{Y}}}={\rm fft}\left({\underline{\mathbf{Y}},[],3}\right)$;
Output : Moore-Penrose pseudoinverse $\underline{\mathbf{X}}^{{\dagger}}\in\mathbb{R}^{I_{2}\times I_{1}\times I_{3}}$
@@ 3       $\widehat{\underline{\mathbf{C}}}\left({:,:,i}\right)={\rm pinv}\,(\widehat{{\bf X}}(:,:,i))$;
Algorithm 2 Fast Moore-Penrose pseudoinverse computation of the tensor $\underline{\bf X}$
@@ 3 Matrix CUR (MCUR) approximation of a single matrix and matrix pairs
The MCUR approximation of matrix pairs (matrices with the same number of columns) as the extension of the classical MCUR approximations were presented in [2]. The key idea of this generalization is to select columns from two matrices that share the same column indices in order to extract the features of one dataset that are most pertinent to the other. Here, the common factor matrix produced by the GSVD of the two matrices is used to sample the columns and rows using the Discrete Empirical Interpolation Method (DEIM) technique. The MCUR approximation of triplet matrices was also presented in [3] , with the goal of identifying pertinent characteristics of one data matrix in relation to two other matrices. The authors in this instance apply the RSVD to produce two common factor matrices, to which the DEIM is then applied to sample the column and row indices. We first discuss the DEIM extension to tensors based on the t-product in the next section before using it to extend the TCUR technique to tensor pairs and tensor triplets in Sections 5 and 6, respectively.
@@ 4 Tensor Discrete Empirical Interpolation Method (TDEIM) for lateral/horizontal slice sampling
We have recently generalized the DEIM method to the tensor case based on the t-product [20] that we call tensor DEIM (TDEIM). It has been experimentally shown that the TDEIM achieved the best sampling accuracy compared to the existing sampling techniques, such as top tubal leverage score sampling and uniform sampling. The TDEIM to select horizontal slices is presented in Algorithm 4. A key concept in the TDEIM is the interpolatory projector, as we now define it. For a given set of indices ${\bf s}\in\mathbb{R}^{R}$, the full tubal-rank tensor [^1] A tensor with linear independent lateral slices, for example, the tensor $\underline{\bf U}$ obtained from the t-SVD can be used. $\underline{\bf U}\in\mathbb{R}^{I_{1}\times R\times I_{3}}$ and considering $\underline{\bf S}=\underline{\bf I}(:,{\bf s},:)\in\mathbb{R}^{I_{1}\times R\times I_{3}}$, as the selection tensor where $\underline{\bf I}\in\mathbb{R}^{R\times R\times I_{3}}$ is an identity tensor, we build the tensorial oblique projection operator defined as follows
@@
$s_{j}=\arg\max_{1\leq i\leq I_{1}}\|\underline{\bf R}(i,j,:)\|,$
is set as our new index $s_{j}$. This will be the new sampled horizontal slice’s index. The same procedure can be used to select lateral slices where the tensor $\underline{\bf V}$ obtained from the t-SVD of the tensor $\underline{\bf X}$ should be used.
@@
that will be used in the next theorem. Here, $\underline{\bf V}\in\mathbb{R}^{I_{2}\times R\times I_{3}}$ is a bais tensor for the subspace of horizontal slices of the original data tensor, $\underline{\bf Q}=\underline{\bf I}(:,{\bf s}_{j-1},:)$ is a tensor of some sampled lateral slices of the identity tensor specified by the index set ${\bf s}_{j-1}=\{s_{1},s_{2},\ldots,s_{j-1}\},\,\widehat{\underline{\bf Q}}={\rm fft}(\underline{\bf Q},[],3),\,\widehat{\underline{\bf V}}={\rm fft}(\underline{\bf V},[],3),\,\widehat{\underline{\bf Q}}_{i}=\widehat{\underline{\bf Q}}(:,:,i),$ and $\widehat{\underline{\bf V}}_{i}=\widehat{\underline{\bf V}}(:,:,i)$.
The next theorem provides the upper error bound of the TCUR approximation obtained by the TDEIM sampling method.
@@
[20] Suppose $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $1\leq R<\min(I_{1},I_{2})$. Assume that horizontal slice and lateral slice indices ${\bf p}$ and ${\bf q}$ give full tubal rank tensors $\underline{\bf C}=\underline{\bf X}(:,{\bf q},:)=\underline{\bf X}*\underline{\mathcal{Q}}$ and $\underline{\bf R}=\underline{\bf X}({\bf p},:,:)={\underline{\mathcal{P}}}*\underline{\bf X}$, where $\underline{\mathcal{P}}$ and $\underline{\mathcal{Q}}$ are tensorial interpolation projectors for horizontal and lateral slice sampling [^2] The tensorial interpolation projector for the lateral slice sampling is defined as $\underline{\mathcal{Q}}=\underline{{\bf V}}*(\underline{{\bf Q}}^{T}*\underline{{\bf V}})^{-1}*\underline{{\bf Q}}^{T}$, where $\underline{{\bf V}}$ is a basis tensor and $\underline{{\bf Q}}$ is the collection tensor., respectively with corresponding finite error constants $\tilde{\eta}_{p},\,\tilde{\eta}_{q}$ defined in (8) and set $\underline{\bf U}=\underline{{\bf C}}^{{\dagger}}*\underline{\bf X}*\underline{\bf R}^{{\dagger}}$. Then
@@ 5 Tensor CUR (TCUR) approximation for tensor pairs based on the t-product
Inspired by the success of the MCUR approximation, several generalizations of these approaches to tensors have been proposed. For example, see [34] for the Tucker model, see [35, 36] for the TT decomposition, see [18] for the t-SVD, and see [37] for the TC/TR decomposition. In this note, we focus on the t-SVD case and extend it to a pair of tensors. The intuition behind this generalization is similar to the matrix case. More precisely, the indices of selected lateral slices are the same for two data tensors (see Figure 1 for a graphical illustration). The formal definition of the GTCUR for tensor pairs is presented below.
Figure 1: Visualization of the generalized tensor CUR (GTCUR) approximation for tensor pairs $(\underline{\bf X},\underline{\bf Y})$. The indices of the selected lateral slices of the tensors $\underline{\bf X}$ and $\underline{\bf Y}$ are identical while this is not necessarily true for the horizontal slices.
@@
Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\underline{\bf Y}\in\mathbb{R}^{I_{4}\times I_{2}\times I_{3}}$ be of full tubal rank, with $I_{1}\geq I_{2}$ and $I_{4}\geq I_{2}$. A GTCUR decomposition of the tensor pairs $(\underline{\bf X},\underline{\bf Y})$ with the tubal rank $R,$ is defined as follows
@@
where $\underline{\bf S}_{1}\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\underline{\bf S}_{2}\in\mathbb{R}^{I_{4}\times R\times I_{3}}$ and $\underline{\bf P}\in\mathbb{R}^{I_{2}\times R\times I_{3}}$ are index the selection tensors ($R<I_{2}$).
Assume that ${\bf s}_{1},$ and ${\bf s}_{2}$ are the horizontal slice indices of the tensors $\underline{\bf X}$ and $\underline{\bf Y}$, respectively. Also, let ${\bf p}$ be the common lateral slice indices for the both tensors $\underline{\bf X}$ and $\underline{\bf Y}$. Then, formulations (10)-(11) can be equavalently written as
@@
The GTCUR for tensor pairs indeed computes the TCUR approximations of two given data tensors $\underline{\bf X}$ and $\underline{\bf Y}$ and is motivated by the generalized t-SVD (GTSVD). For given data tensors $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\underline{\bf Y}\in\mathbb{R}^{I_{4}\times I_{2}\times I_{3}}$, the GTSVD guarantees the following decompositions [17]